OpenStudy (anonymous):

solve: 16x^2+24x+9=-16

OpenStudy (mertsj):

Add 16 to both sides first.

OpenStudy (mertsj):

Add 16 to both sides first.

OpenStudy (hoblos):

16x^2+24x+9=-16 (4x)^2 + 2(4x)(3) + 3^2 = -16 (4x+3)^2 = -16 which is impossible since a squared number is always positive

OpenStudy (amistre64):

its not impossible; it just doesnt have "real" roots

OpenStudy (anonymous):

1) 16x^2+24x=-25 2) (4x+3)(4x+3)=-25+9 3) Answer:(4x+3)(4x+3)+16